# The equilibrium properties of obvious strategy profiles

### v1

@@ 2.2 Large finite-player games
In this subsection, we introduce a class of large finite-player games. Let a finite probability space $(I^{n},\mathcal{I}^{n},\lambda^{n})$ denote the set of players. Here we assume that $|I^{n}|=n$ and $\mathcal{I}^{n}$ consists of all the subsets of $I^{n}$ (i.e., the power set of $I^{n}$). For simplicity, let $A$ be the common action space for all the players in this game.
Moreover, each player’s payoff function depends on her own choice and the probability distribution on $A$ that is induced from the action profile (i.e., the societal summary). Clearly, the set of such action distributions is a subset of $\mathcal{M}(A)$ that is denoted by
$D^{n}=\Bigl{\{}\tau\in\mathcal{M}(A)\;\Big{|}\;\tau=\sum\limits_{j\in I^{n}}\lambda^{n}(j)\delta_{a_{j}}\text{ where }a_{j}\in A\text{ for all }j\in I^{n}\Bigr{\}},$
where $\delta_{a}$ denotes the Dirac probability measure that assigns probability one to $\{a\}$, for all $a\in A$. Player $i$’s payoff function is then given by a bounded continuous function $u_{i}^{n}\colon A\times\mathcal{M}(A)\to\mathbb{R}$, clearly, $u_{i}^{n}\in\mathcal{U}_{A}$. Thus, a large finite-player game $G^{n}$ can be viewed as a mapping from $I^{n}$ to $\mathcal{U}_{A}$ such that $G^{n}(i)=u_{i}^{n}$ for all $i\in I^{n}$.
In this finite-player game, a pure strategy profile $f^{n}$ is a mapping from $(I^{n},\mathcal{I}^{n},\lambda^{n})$ to $A$. Hence, given a pure strategy profile $f^{n}$, the payoff function for player $i$ is
$u_{i}^{n}(f^{n})=u_{i}^{n}\Bigl{(}f^{n}(i),\sum\limits_{j\in I^{n}}\lambda^{n}(j)\delta_{f^{n}(j)}\Bigr{)},$
here we slightly abuse the notation $u_{i}^{n}(f^{n})$ to denote player $i$’s payoff given the strategy profile $f^{n}$.
Similarly, a randomized strategy is a probability distribution $\mu\in\mathcal{M}(A)$. A randomized strategy profile $g^{n}$ is a mapping from $(I^{n},\mathcal{I}^{n},\lambda^{n})$ to $\mathcal{M}(A)$. Thus, given a randomized strategy profile $g^{n}$, player $i$’s (expected) payoff is
$u_{i}^{n}(g^{n})=\int_{A^{n}}u_{i}^{n}\Bigl{(}a_{i},\sum\limits_{j\in I^{n}}\lambda^{n}(j)\delta_{a_{j}}\Bigr{)}\mathbin{\mathop{\otimes}\limits_{j\in I^{n}}}g^{n}(j,{\rm{d}}a_{j}),$
where $\mathbin{\mathop{\otimes}\limits_{j\in I^{n}}}g^{n}(j,{\rm{d}}a_{j})$ is the product probability measure on the product space $A^{n}$. The societal summary induced by $g^{n}$ is $s(g^{n})=\int_{I^{n}}g^{n}(i)\mathrm{d}\lambda^{n}(i)$. Moreover, a randomized strategy profile $g^{n}$ is said to be symmetric if for any two players $i$ and $i^{\prime}$, $g^{n}(i)=g^{n}(i^{\prime})$ whenever $G^{n}(i)=G^{n}(i^{\prime})$. Finally, we state the definitions of randomized strategy Nash equilibrium and $\varepsilon$-Nash equilibrium as follows.
@@
A randomized strategy profile $g^{n}\colon I^{n}\rightarrow\mathcal{M}(A)$ is said to be a randomized strategy Nash equilibrium if for all $i\in I^{n}$,
@@
where $(\mu,g^{n}_{-i})$ represents the randomized strategy profile such that player $i$ plays the randomized strategy $\mu$, and player $j$ plays the randomized strategy $g^{n}(j)$ for all $j\in I^{n}\backslash\{i\}$.
@@
For any $\varepsilon\geq 0$, a randomized strategy profile $g^{n}\colon I^{n}\to\mathcal{M}(A)$ is said to be an $\varepsilon$-Nash equilibrium if there exists a subset of players $I^{n}_{\varepsilon}\subseteq I^{n}$ such that $\lambda^{n}(I^{n}_{\varepsilon})\geq 1-\varepsilon$ and for all $i\in I^{n}_{\varepsilon}$,
@@ 3 Main results
In this section, we establish the main results.. Let $G^{n}$ be a large finite-player game with $n$ players. The players in this game use a large game $G$ with a continuum of players as an approximation to simplify the equilibrium analysis. We assume that the players’ characteristic information (i.e., payoff functions) from the finite-player game $G^{n}$ is included in the imagined large game $G$; specifically, we have $G^{n}(I^{n})\subset\operatorname{supp}\lambda G^{-1}$. [^9] Since $G$ is a measurable mapping from the player space $(I,\mathcal{I},\lambda)$ to the characteristic space $\mathcal{U}_{A}$, $\lambda G^{-1}$ is the induced measure on $\mathcal{U}_{A}$. The support of $\lambda G^{-1}$, denoted as $\mathrm{supp}\lambda G^{-1}$, is the smallest closed set $B\in\mathcal{B}(\mathcal{U}_{A})$ such that $\lambda G^{-1}(B)=1$. As we will demonstrate, this assumption allows the players of the large finite-player game $G^{n}$ to implement an equilibrium outcome of the large game $G$ directly.
Suppose that $\tau$ is an action distribution induced by some Nash equilibrium $g$ of large game $G$. This equilibrium action distribution $\tau$ is considered a desirable social choice outcome for players in the large finite game $G^{n}$. As discussed in Section 1, each player may have multiple optimal choices when trying to implement the social choice outcome $\tau$. To address this coordination issue, we adopt a process to symmetrize the equilibrium action distribution. Specifically, we can find a symmetric Nash equilibrium $\widetilde{g}$ of the large game $G$ such that the societal summary of $\widetilde{g}$ also equals $\tau$, i.e., $s(\widetilde{g})=\tau$. This symmetrization result is possible due to the existence of an auxiliary mapping defined as follows.
@@
Given an equilibrium action distribution $\tau$, the auxiliary mapping associated with $\tau$ is a mapping $\overline{g}\colon\mathcal{U}_{A}\to\mathcal{M}(A)$ such that the composition mapping $\overline{g}\circ G\colon I\to\mathcal{M}(A)$ (denoted by $\widetilde{g}$) is a Nash equilibrium of $G$, and satisfies $s(\widetilde{g})=\tau$.
@@
Figure 1: Auxiliary mapping
Notice that $\overline{g}$ assigns a strategy to each characteristic. Thus, in the strategy profile $\widetilde{g}=\overline{g}\circ G$, players with the same characteristic choose the same strategy. This implies that $\widetilde{g}$ constitutes a symmetric Nash equilibrium with the aggregate action distribution $\tau$.
@@
For any equilibrium action distribution $\tau$ of a large game $G$, the auxiliary mapping $\overline{g}$ always exists.
The proof of Lemma 1 is provided in Subsection 5.2.101010Furthermore, it can be shown that given any large game and equilibrium action distribution, the auxiliary mapping is unique. This proof serves as the foundation for the symmetrization process. Clearly, playing a symmetric strategy profile $\widetilde{g}$ avoids coordination issues, as players only need to act according to their characteristics. Since the aggregate action distribution, or societal summary, of $\widetilde{g}$ also equals $\tau$, players can simply focus on the symmetric equilibrium $\widetilde{g}$ to implement $\tau$.
@@
Each player $i$ in the large finite-player game $G^{n}$ has an imagined partner $i^{\prime}$ in the large game $G$ that shares the same characteristics, i.e., $G^{n}(i)=G(i^{\prime})$. [^11] Here we apply the assumption introduced at the beginning of this subsection: $G^{n}(I^{n})\subset\operatorname{supp}\lambda G^{-1}$. A straightforward method for players in $G^{n}$ to implement $\tau$ is for each player to adopt the strategy employed by her imagined partner in $G$, represented as $\widetilde{g}(i^{\prime})=\overline{g}\circ G(i^{\prime})=\overline{g}\circ G^{n}(i)$. This resulting strategy profile in $G^{n}$ is referred to as the obvious strategy profile, which is formally defined as follows.
Definition 6 (Obvious strategy profile).
@@
Given a large finite-player game $G^{n}$, a large game $G$ approximating $G^{n}$, and an auxiliary mapping $\overline{g}$, an obvious strategy profile of $G^{n}$ is defined as $g^{n}=\overline{g}\circ G^{n}$.
@@
Figure 2: Obvious stratgy profile
Notice that the obvious strategy profile $g^{n}$ is symmetric, as players with the same characteristics select the same strategy determined by $\overline{g}$. Each player in this strategy profile $g^{n}$ chooses the strategy that corresponds to their imagined partner in the strategy profile $\widetilde{g}$ of the large game $G$. We now revisit the motivating example in Section 1 to illustrate $\widetilde{g}$, $\overline{g}$, and $g^{n}$. In that example, $\widetilde{g}(i)=\tfrac{1}{2}\delta_{B}+\tfrac{1}{2}\delta_{H}$ for each player $i$ in the large Beach vs. Hotel game. Since players share a common payoff function $u$ in that game, the characteristic space $G(I)$ is the singleton $u$, and thus $\overline{g}(u)=\tfrac{1}{2}\delta_{B}+\tfrac{1}{2}\delta_{H}$. Therefore, for each finite Beach vs. Hotel game $G^{n}$ with $n$ players, the obvious strategy profile is $g^{n}(i)=\tfrac{1}{2}\delta_{B}+\tfrac{1}{2}\delta_{H}$ for each player $i\in I^{n}$.
Our main theorem examines the equilibrium property of the obvious strategy profile $g^{n}$. According to our analysis in Section 1, $g^{n}$ is not generally an exact Nash equilibrium; therefore, it can be viewed as an $\varepsilon_{n}$-Nash equilibrium for some $\varepsilon_{n}\geq 0$. As $n$ varies, we obtain a sequence of symmetric obvious strategy profiles $\{g^{n}\}_{n\in\mathbb{Z}_{+}}$ corresponding to the sequence of finite-player games $\{G^{n}\}_{n\in\mathbb{Z}_{+}}$. This leads to a natural question: does the sequence $\{g^{n}\}_{n\in\mathbb{Z}_{+}}$ converge? More specifically, does $\varepsilon_{n}\to 0$ as the games $\{G^{n}\}_{n\in\mathbb{Z}_{+}}$ converge to $G$? To address this question, we first introduce the following concept regarding the convergence of games.
Definition 7.
A sequence of finite-player games $\{G^{n}\}_{n\in\mathbb{Z}_{+}}$ converges to a large game $G$ if the distributions of characteristics of $\{G^{n}\}_{n\in\mathbb{Z}_{+}}$ converge to the distribution of characteristics of $G$. Specifically, this means that $\{\lambda^{n}(G^{n})^{-1}\}_{n\in\mathbb{Z}_{+}}$ converges weakly to $\lambda G^{-1}$.
Now we are ready to present the first main result. In Theorem 1 below, we show that if the auxiliary mapping $\overline{g}$ is almost everywhere continuous on the subset $\mathrm{supp}\lambda G^{-1}\subset\mathcal{U}_{A}$, then the sequence of obvious strategy profiles $\{g^{n}\}_{n\in\mathbb{Z}_{+}}$ is a convergent sequence of approximate Nash equilibria.
@@
Given a sequence of finite game $\{G^{n}\}_{n\in\mathbb{Z}_{+}}$ that converges to a large game $G$, let $\overline{g}$ be an auxiliary mapping of $G$. Define $\{g^{n}\}_{n\in\mathbb{Z}_{+}}$ as the sequence of obvious strategy profiles induced by $\overline{g}$. If $\overline{g}$ is almost everywhere continuous on $\mathrm{supp}\lambda G^{-1}$, then there exists a sequence of real numbers $\{\varepsilon_{n}\}_{n\in\mathbb{Z}_{+}}$ such that each $g^{n}$ is an $\varepsilon_{n}$-Nash equilibrium, and $\varepsilon_{n}\to 0$ as $n$ approaches infinity.
The proof of Theorem 1, which is collected in Section 5, comprises two main components and presents two technical contributions. In Subsection 5.1, as the first technical contribution, we establish a theory on the convergence of realized society summaries of randomized strategy profiles. This convergence result is crucial in our proof since an obvious strategy profile $g^{n}$ generally involves randomized strategies. In Subsection 5.3, we make the second technical contribution by decomposing the remaining proof into five steps. We then establish a series of auxiliary inequalities pivotal to the proof. Finally, we finish the proof by using auxiliary inequalities together with the continuity condition on $\overline{g}$.
Notice that the continuity requirement of $\overline{g}$ in Theorem 1 is crucial. Below we present an example showing that if the continuity assumption of $\overline{g}$ is not satisfied, then a sequence of obvious strategy profiles induced by $\overline{g}$ may not be a convergent sequence of approximate symmetric equilibria (i.e., $\{\varepsilon_{n}\}_{n\in\mathbb{Z}_{+}}$ does not converge to 00 as $n$ approaches infinity).

### v2

@@ 2.2 Large finite-player games
In this subsection, we introduce a class of large finite-player games. Let $(I_{n},\mathcal{I}_{n},\lambda_{n})$ be a finite probability space representing the set of players. We assume that $|I_{n}|=n$, that $\mathcal{I}_{n}$ is the power set of $I_{n}$, and that $\lambda_{n}(i)=\tfrac{1}{n}$ for each $i\in I_{n}$. [^9] The assumption that each player has equal weight is not essential, but we adopt it to simplify the model. For simplicity, we assume that all players share a common action space $A$.
Similarly, each player’s payoff function depends on her own choice and the probability distribution on $A$ that is induced from the action profile (i.e., the societal summary). Clearly, the set of such action distributions is a subset of $\mathcal{M}(A)$ that is denoted by
$D^{n}=\Bigl{\{}\tau\in\mathcal{M}(A)\;\Big{|}\;\tau=\sum\limits_{j\in I_{n}}\lambda_{n}(j)\delta_{a_{j}}\text{ where }a_{j}\in A\text{ for all }j\in I_{n}\Bigr{\}},$
Player $i$’s payoff function is then given by a bounded continuous function $u_{i}^{n}\colon A\times\mathcal{M}(A)\to\mathbb{R}$, clearly, $u_{i}^{n}\in\mathcal{U}_{A}$. Thus, a large finite-player game $G_{n}$ can be represented as a mapping from $I_{n}$ to $\mathcal{U}_{A}$, with $G_{n}(i)=u_{i}^{n}$ for all $i\in I_{n}$.
In this finite-player game, a pure strategy profile $f^{n}$ is a mapping from $(I_{n},\mathcal{I}_{n},\lambda_{n})$ to $A$. Hence, given a pure strategy profile $f^{n}$, the payoff function for player $i$ is
$u_{i}^{n}(f^{n})=u_{i}^{n}\Bigl{(}f^{n}(i),\sum\limits_{j\in I_{n}}\lambda_{n}(j)\delta_{f^{n}(j)}\Bigr{)},$
here we slightly abuse the notation $u_{i}^{n}(f^{n})$ to denote player $i$’s payoff given the strategy profile $f^{n}$. Similarly, a randomized strategy is a probability distribution $\mu\in\mathcal{M}(A)$. A randomized strategy profile $g^{n}$ is a mapping from $(I_{n},\mathcal{I}_{n},\lambda_{n})$ to $\mathcal{M}(A)$. Thus, given a randomized strategy profile $g^{n}$, player $i$’s (expected) payoff is given by
$u_{i}^{n}(g^{n})=\int_{A^{n}}u_{i}^{n}\Bigl{(}a_{i},\sum\limits_{j\in I_{n}}\lambda_{n}(j)\delta_{a_{j}}\Bigr{)}\mathbin{\mathop{\otimes}\limits_{j\in I_{n}}}g^{n}(j,{\rm{d}}a_{j}),$
where $\mathbin{\mathop{\otimes}\limits_{j\in I_{n}}}g^{n}(j,{\rm{d}}a_{j})$ is the product probability measure on the product space $A^{n}$. The societal summary induced by $g^{n}$ is $s(g^{n})=\int_{I_{n}}g^{n}(i)\mathrm{d}\lambda_{n}(i)$. Moreover, a randomized strategy profile $g^{n}$ is said to be symmetric if for any two players $i$ and $i^{\prime}$, $g^{n}(i)=g^{n}(i^{\prime})$ whenever $G_{n}(i)=G_{n}(i^{\prime})$. Finally, we state the definitions of randomized strategy Nash equilibrium and $\varepsilon$-Nash equilibrium as follows.
@@
A randomized strategy profile $g^{n}\colon I_{n}\rightarrow\mathcal{M}(A)$ is said to be a randomized strategy Nash equilibrium if for all $i\in I_{n}$,
@@
where $(\mu,g^{n}_{-i})$ represents the randomized strategy profile such that player $i$ plays the randomized strategy $\mu$, and player $j$ plays the randomized strategy $g^{n}(j)$ for all $j\in I_{n}\backslash\{i\}$.
@@
For any $\varepsilon\geq 0$, a randomized strategy profile $g^{n}\colon I_{n}\to\mathcal{M}(A)$ is said to be an $\varepsilon$-Nash equilibrium if there exists a subset of players $I_{n}^{\varepsilon}\subseteq I_{n}$ such that $\lambda_{n}(I_{n}^{\varepsilon})\geq 1-\varepsilon$ and for all $i\in I_{n}^{\varepsilon}$,
@@ 3 Main results
In this section, we present the main results. Let $G_{n}$ be a large finite-player game with $n$ players. To simplify the equilibrium analysis, we assume that the players in $G_{n}$ use a large game GG—featuring a continuum of players—as an approximation. Specifically, we assume that the characteristic information (i.e., the payoff functions) of the players in $G_{n}$ is contained within that of the large game $G$. Formally, this means $G_{n}(I_{n})\subset\mathrm{supp}\,\lambda G^{-1}$. [^10] Since $G$ is a measurable mapping from the player space $(I,\mathcal{I},\lambda)$ to the characteristic space $\mathcal{U}_{A}$, the measure $\lambda G^{-1}$ represents the distribution of characteristics among players. Its support, denoted by $\mathrm{supp}\,\lambda G^{-1}$, is the smallest closed subset of $\mathcal{U}_{A}$ that has full measure. As we will show, this assumption ensures that the players in the finite-player game $G_{n}$ can directly implement an equilibrium outcome of the large game $G$.
Suppose that $\tau^{*}$ is the societal summary (i.e., the action distribution) induced by a Nash equilibrium $g$ of the large game $G$. This equilibrium action distribution $\tau^{*}$ is considered a desirable social choice outcome for the players in the finite-player game $G_{n}$. As discussed in Section 1, when attempting to implement the outcome $\tau^{*}$, each player may face multiple optimal responses, creating a coordination problem. To resolve this issue, we adopt a process to symmetrize the equilibrium action distribution. Specifically, there exists a symmetric Nash equilibrium $\widetilde{g}$ of the large game $G$ such that its societal summary also equals $\tau^{*}$; that is, $s(\widetilde{g})=\tau^{*}$. This symmetrization result relies on the existence of an auxiliary mapping, defined as follows.
@@
Given an equilibrium action distribution $\tau^{*}$, the auxiliary mapping associated with $\tau^{*}$ is a function $\overline{g}\colon\mathcal{U}_{A}\to\mathcal{M}(A)$ such that the composition $\widetilde{g}\mathrel{\mathop{\ordinarycolon}}=\overline{g}\circ G\colon I\to\mathcal{M}(A)$ is a Nash equilibrium of the large game $G$ and satisfies $s(\widetilde{g})=\tau^{*}$.
@@
Figure 1: Construction of a symmetric equilibrium via the auxiliary mapping
Note that the auxiliary mapping $\overline{g}$ assigns a strategy to each possible player characteristic. As a result, in the composed profile $\widetilde{g}=\overline{g}\circ G$, players who share the same characteristic adopt the same strategy. Lemma 1 below guarantees the existence of such a symmetric equilibrium $\widetilde{g}$ whose societal summary coincides with $\tau^{*}$.
@@
For any equilibrium action distribution $\tau^{*}$ of a large game $G$, the auxiliary mapping $\overline{g}$ always exists.
Lemma 1 follows as a corollary of Sun et al. (2020, Theorem 2(ii)), which relies on the technique of disintegration to establish the existence of $\overline{g}$. Since our model focuses on games with a common action set, we provide a shorter and more direct proof in Subsection 5.2. Notably, playing a symmetric strategy profile $\widetilde{g}$ resolves coordination issues, as players need only respond based on their characteristics. Moreover, since the societal summary of $\widetilde{g}$ coincides with $\tau^{*}$, players can coordinate on the symmetric equilibrium $\widetilde{g}$ to implement the desired equilibrium outcome $\tau^{*}$.
@@
Each player $i$ in the finite-player game $G_{n}$ is associated with an imagined partner $i^{\prime}$ in the large game $G$ who shares the same characteristic; that is, $G_{n}(i)=G(i^{\prime})$. [^11] This relies on the assumption introduced at the beginning of this section: $G_{n}(I_{n})\subset\mathrm{supp}\,\lambda G^{-1}$. A natural way for players in $G_{n}$ to implement the desired societal summary $\tau^{*}$ is to imitate the strategy used by their imagined partners in $G$. Specifically, player $i$ adopts the strategy $\widetilde{g}(i^{\prime})=\overline{g}\circ G(i^{\prime})=\overline{g}\circ G_{n}(i)$. The resulting strategy profile in $G_{n}$ is referred to as the direct strategy profile, formally defined below.
Definition 6 (Direct strategy profile).
@@
Given a large finite-player game $G_{n}$, a large game $G$, and an associated auxiliary mapping $\overline{g}$, the direct strategy profile of $G_{n}$ is defined by $g^{n}=\overline{g}\circ G_{n}$.
$I_{n}$$\mathcal{M}(A)$$\mathcal{U}_{A}$$g^{n}$$G_{n}$$\overline{g}$
Figure 2: Direct strategy profile
Observe that the direct strategy profile $g^{n}$ is symmetric, as players with the same characteristic choose the same strategy, determined by the auxiliary mapping $\overline{g}$. In this strategy profile, each player adopts the strategy that her imagined partner would use in the symmetric equilibrium $\widetilde{g}$ of the large game $G$.
We now revisit the motivating example from Section 1 to illustrate the roles of $\widetilde{g}$, $\overline{g}$, and $g^{n}$. In that example, $\widetilde{g}(i)=\tfrac{1}{3}\delta_{a}+\tfrac{2}{3}\delta_{b}$ for every player $i$ in the large routing game. Since all players share the same payoff function $u$, the image $G(I)$ is the singleton $\{u\}$, and thus $\overline{g}(u)=\tfrac{1}{3}\delta_{a}+\tfrac{2}{3}\delta_{b}$. Consequently, for any finite routing game $G_{n}$ with $n$ drivers, the corresponding direct strategy profile satisfies $g^{n}(i)=\tfrac{1}{3}\delta_{a}+\tfrac{2}{3}\delta_{b}$ for each driver $i\in I_{n}$.
Our main theorems examine the equilibrium properties of the direct strategy profile $g^{n}$. As discussed in Section 1, $g^{n}$ is generally not an exact Nash equilibrium; instead, it can be regarded as an $\varepsilon_{n}$-Nash equilibrium for some $\varepsilon_{n}\geq 0$. Varying $n$, we obtain a sequence of direct strategy profiles $\{g^{n}\}_{n\in\mathbb{Z}_{+}}$ corresponding to the sequence of finite-player games $\{G_{n}\}_{n\in\mathbb{Z}_{+}}$. This raises a natural question: does the sequence $\{g^{n}\}_{n\in\mathbb{Z}_{+}}$ converge? More specifically, does $\varepsilon_{n}$ tend to zero as the games $\{G_{n}\}_{n\in\mathbb{Z}_{+}}$ converge to $G$? To answer this, we first introduce the following concept regarding the convergence of games.
Definition 7 (Convergence of games).
A sequence of finite-player games $\{G_{n}\}_{n\in\mathbb{Z}_{+}}$ converges to a large game $G$ if the corresponding distributions of player characteristics converge weakly; that is,
$\lambda_{n}\circ G_{n}^{-1}\xrightarrow{w}\lambda\circ G^{-1}\quad\text{on }\mathcal{U}_{A}.$
Definition 7 formalizes a widely used notion of convergence in the study of large games: the weak convergence of the distributions of player characteristics. Under this notion, the sequence of characteristic distributions in the finite-player games $\{G_{n}\}_{n\in\mathbb{Z}_{+}}$ converges to the limiting distribution induced by the large game $G$. The game-theoretic interpretation is that, as the number of players grows, individual heterogeneity becomes negligible, and the aggregate structure of the game stabilizes. This convergence justifies using large games as approximations for analyzing strategic behavior in large but finite populations.
We are now ready to present our first main result. Theorem 1 shows that if the auxiliary mapping $\overline{g}$ is almost everywhere continuous on the support of the characteristic distribution (i.e., $\mathrm{supp}\,\lambda G^{-1}$), then the sequence of direct strategy profiles $\{g^{n}\}_{n\in\mathbb{Z}_{+}}$—each constructed by applying $\overline{g}$ to the characteristics in $G_{n}$—forms a convergent sequence of approximate Nash equilibria.
Theorem 1.
Let $\{G_{n}\}_{n\in\mathbb{Z}_{+}}$ be a sequence of finite-player games that converges to a large game $G$. Let $\overline{g}$ be an auxiliary mapping associated with $G$, and define $\{g^{n}\}_{n\in\mathbb{Z}_{+}}$ as the sequence of direct strategy profiles induced by $\overline{g}$. If $\overline{g}$ is almost everywhere continuous on $\mathrm{supp}\,\lambda G^{-1}$, then there exists a sequence $\{\varepsilon_{n}\}_{n\in\mathbb{Z}_{+}}$ with $\varepsilon_{n}\to 0$ such that each $g^{n}$ is an $\varepsilon_{n}$-Nash equilibrium of $G_{n}$.
Example 1 in Subsection 4.1 below illustrates that the continuity condition imposed on the auxiliary mapping $\overline{g}$ is necessary for the conclusion of Theorem 1 to hold. Nonetheless, this requirement is relatively mild and satisfied by many classes of large games studied in the literature. A particularly important case occurs when the large game features only finitely many player characteristics—that is, when $G(I)$ is a finite subset of $\mathcal{U}_{A}$. In such settings, $\overline{g}$ is trivially almost everywhere continuous, since any function defined on a finite set is continuous on its support. Consequently, Theorem 1 applies directly, and as shown in Corollary 1, the sequence of direct strategy profiles constructed from the corresponding finite-player games forms a convergent sequence of approximate symmetric equilibria.
Corollary 1.
Let $G\colon(I,\mathcal{I},\lambda)\to\mathcal{U}_{A}$ be a large game such that $G(I)$ is a finite subset of $\mathcal{U}_{A}$, and let $\{G_{n}\}_{n\in\mathbb{Z}_{+}}$ be a sequence of finite games converging to $G$. Suppose $\{g^{n}\}_{n\in\mathbb{Z}_{+}}$ is the sequence of direct strategy profiles induced by the auxiliary mapping $\overline{g}$. Then, there exists a sequence $\{\varepsilon_{n}\}_{n\in\mathbb{Z}_{+}}$ with $\varepsilon_{n}\to 0$ such that each $g^{n}$ is an $\varepsilon_{n}$-Nash equilibrium of $G_{n}$.
This result has practical significance. In many real-world applications, the number of distinct player characteristics is limited, effectively making the characteristic space finite. Therefore, Theorem 1 together with Corollary 1 imply that, in large finite-player games, it is asymptotically optimal for players to adopt the strategies played by their imagined partners in the corresponding large game.
Remark 1.
Theorem 1 can be extended to games in which players may have different feasible action spaces—i.e., each player’s characteristic consists of both a feasible action set and a payoff function. In this more general setting, the notions of auxiliary mapping and direct strategy profile can be defined in a similar manner. Moreover, if the auxiliary mapping is almost everywhere continuous on $\mathrm{supp}\,\lambda G^{-1}$, then the sequence of induced direct strategy profiles converges.
Remark 2.
Beyond the setting of games with finitely many characteristics, the continuity condition on $\bar{g}$ is also satisfied in many games with infinitely many characteristics. For instance, consider the case where both the characteristic mapping $G\colon I\to\mathcal{U}_{A}$ and the strategy profile $\widetilde{g}\colon I\to\mathcal{M}(A)$ are continuous. [^12] Here, $\mathcal{U}_{A}$ is endowed with the sup-norm topology, $I$ is equipped with the topology induced by $G$, and $\mathcal{M}(A)$ is endowed with the weak topology. Then, as illustrated in Figure 1, the auxiliary mapping $\overline{g}$ is also continuous. For related discussions on continuous Nash equilibria in games with continuous player spaces, see Kim (1997). [^13] Kim (1997) adopts the argmax topology on the characteristic space $\mathcal{U}_{A}$, which differs from the sup-norm topology used here.
We conclude this section by presenting our second main result, which examines the equilibrium property of direct strategy profiles after the resolution of uncertainty. Given a sequence $\{g^{n}\}_{n\in\mathbb{Z}_{+}}$, where each $g^{n}$ is a randomized strategy profile in the finite-player game $G_{n}$, standard probability theory ensures the existence of a probability space $(\Omega,\Sigma,\mathbb{P})$ and a sequence of measurable mappings $\{x^{n}\colon I_{n}\times\Omega\to A\}_{n\in\mathbb{Z}_{+}}$ such that:
  • (i)
    For each $i\in I_{n}$ and $n\in\mathbb{Z}_{+}$, the distribution of the random variable $x^{n}_{i}\mathrel{\mathop{\ordinarycolon}}=x^{n}(i,\cdot)$ coincides with $g^{n}(i)$;
  • (ii)
    For each $n\in\mathbb{Z}_{+}$, the family of random variables $\{x^{n}_{i}\}_{i\in I_{n}}$ are pairwise independent.
For each $\omega\in\Omega$, the mapping $x^{n}(\cdot,\omega)\colon I_{n}\to A$ defines a pure strategy profile, called a realization of $g^{n}$.
For each $n\in\mathbb{Z}_{+}$ and $\varepsilon\geq 0$, we consider the subset $\Omega^{n}_{\varepsilon}\subseteq\Omega$ consisting of realizations that are $\varepsilon$-Nash equilibria in pure strategies:
@@
We are now ready to state the second main result, which shows that realizations of the direct strategy profiles form a convergent sequence of approximate pure strategy equilibria with probability approaching one.
@@
Given a sequence of finite-player game $\{G_{n}\}_{n\in\mathbb{Z}_{+}}$ that converges to a large game $G$, let $\overline{g}$ be an auxiliary mapping of $G$. Define $\{g^{n}\}_{n\in\mathbb{Z}_{+}}$ as the sequence of direct strategy profiles induced by $\overline{g}$. If $\overline{g}$ is almost everywhere continuous on $\mathrm{supp}\lambda G^{-1}$, then for any $\varepsilon>0$,
$\lim\limits_{n\to\infty}\mathbb{P}(\Omega^{n}_{\varepsilon})=1.$
As in Theorem 1, the continuity assumption on $\overline{g}$ in Theorem 2 is also essential, as illustrated by Example 1 in Subsection 4.1. Together, Theorems 1 and 2 establish that, for players in a large finite-player game, adopting the direct strategy profile is asymptotically optimal—not only ex ante, in expectation, but also ex post, with high probability.
4 Discussion
This section is divided into three parts. Subsection 4.1 presents a counterexample demonstrating that the continuity condition imposed in Theorems 1 and 2 is essential. Subsection 4.2 investigates direct strategy profiles in pure strategies, showing that, under certain conditions, such strategy profiles exist and form a convergent sequence. Finally, Subsection 4.3 reviews related contributions from the existing literature.
4.1 Failure of convergence without the continuity condition
In this subsection, we demonstrate that the continuity assumption on the auxiliary mapping $\overline{g}$ in Theorems 1 and 2 is essential. We construct a counterexample showing that, if $\overline{g}$ fails to be continuous, then the resulting sequence of direct strategy profiles may fail to form a convergent sequence of approximate symmetric equilibria—specifically, the associated approximation errors $\{\varepsilon_{n}\}_{n\in\mathbb{Z}_{+}}$ may not converge to zero as $n\to\infty$.